# Key-Grid: Unsupervised 3D Keypoints Detection using Grid Heatmap Features

### v1

@@ 1 Introduction
Representing objects through a set of 3D keypoints is one of the most popular and intuitive approaches to compress and comprehend 3D objects [29; 24]. Effectively exhibiting their utility, 3D keypoints have contributed to the success of a number of downstream tasks, including pose estimation, shape registration, object tracking in computer vision [13; 25; 4; 7; 42; 26; 32], as well as various kinds of robotic manipulation tasks [38; 1; 14].
@@
To make the detected keypoints as capable and accessible as possible, the research community is now concentrating on the unsupervised learning of semantically consistent keypoints on 3D point clouds. The implication of semantic consistency is typically twofold: the keypoints should be located at the semantically salient parts of the objects; they are also desired to be aligned within the same category even under large shape variations among diverse 3D object instances. To achieve these objectives, previous works [24; 8; 39; 3; 44] often adopt autoencoder frameworks to facilitate self-supervised training, where the encoders serving as the keypoint predictor are backbone networks [19; 20] generalizable to the shape variation, and the decoders reconstruct the input shape conditioned on the predicted keypoints. Since neural networks are in general better at compression than generation [18], the primary challenge of this pipeline lies in reconstructing the entire point cloud from a few estimated keypoints. Thus, the state-of-the-art (SOTA) detectors [24; 8] lay emphasis on leveraging different priors on 3D structures (e.g., “skeleton” in Skeleton Merger [24], and “cage” in KeypointDeformer [8]) so that the 3D object shapes can be more reasonably approximated by the information from the detected keypoints alone.
While maintaining semantic consistency is already demanding under the shape variations of rigid-body objects (e.g., those in the ShapeNetCoreV2 [2] dataset), it would be even more challenging if deformable objects are also taken into consideration. For instance, when detecting keypoints on the trousers, if one detected keypoint is located at one of the trouser legs, then the keypoint is desired to follow the motion of the trouser leg in the process of the trousers being folded (shown in Figure 1). Note that the shape variation caused by deformation is so dramatic that even the outline of the object shape has been completely altered, indicating a shift in the spatial and geometric structure of the keypoints. Despite the difficulty, with a growing interest in deformable object manipulation [1; 22; 23] in robotics as well as the emergence of large-scale datasets for deformable objects [47; 6] in computer vision, it is of increasing significance to develop a keypoint detector that is equally effectively when faced with deformable objects.
@@ 2 Related works
There have been various approaches proposed for supervised estimation of 3D keypoints using manually annotated keypoints [33; 5; 12; 50; 11; 40]. In contrast to supervised methods, our approach is unsupervised, meaning it does not rely on manually annotated keypoints. Thus, we review methods that adopt an unsupervised approach for estimating 3D keypoints. USIP [10] is the first detector that identifies 3D keypoints in an unsupervised manner by minimizing the chamfer distance between detected keypoints in randomly transformed object pairs. Canonical Capsules [27] is a similar approach that feeds pairs of a randomly translated copies of the same object into a network to detect keypoints. Following USIP, SC3K [49] also uses a random rotation to create two transformed visions of objects and generate the corresponding keypoints by mapping the keypoints of each version back to the original object. Another way to estimate 3D keypoints is proposed by Chen et al. [3]. They encode the point cloud as a set of local feature and input it into a novel structure model to generate the possibility of keypoints. Recently, Fernandez et al. [9] proposed a novel method that distinguishes between Node branches and Pose and coefficient branches to find the optimal keypoints of an object.
@@ 3.3 Decoder: Point Could Reconstruction
As an inverse process of the encoder, the decoder tries to reconstruct the entire point cloud by gradually augmenting finer geometric details in a hierarchical manner. Unlike previous methods [24; 8] where the keypoint-related 3D structures are used only once as the input of the encoder, we propose to integrate increasingly finer grid heatmap features with the layers of the decoder. In addition, the features in the encoder are directly copied and fused into the corresponding layer in the decoder in a U-Net [21] fashion to further assist the reconstruction process.
@@ 3.4 Training Objectives
The common way to evaluate the similarity between a reconstructed point cloud $\mathbf{X}_{r}$ and a target point cloud $\mathbf{X}$ is by employing the Chamfer distance metric [8; 46]. By minimizing the Chamfer distance between them, we can optimize the reconstruction process to closely match the target point cloud, thereby achieving a higher level of similarity between the two point clouds. Thus, we utilize the Chamfer distance between the reconstructed point cloud $\mathbf{X}_{r}$ and the target point cloud $\mathbf{X}$ to approximate this similarity loss $\mathcal{L}_{sim}$.
@@ 4.1 Setup
We use the ShapeNetCoreV2 and the ClothesNet datasets [2; 47] to evaluate the performance of Key-Grid. For the ShapeNetCoreV2 dataset [2], it contains 51,300 rigid-body objects of 55 different categories. In this paper, we only use categories that are manually annotated with semantic correspondence labels in the KeypointNet dataset [40]. For the ClothesNet dataset [47], we take three types of deformations on different type of clothing objects: dropping, dragging, and folding. For each deformation of different garment categories, there are 128 samples during the deformation process. We test the performance of Key-Grid on real-world 3D scans of clothing from the Deep Fashion3D V2 dataset [6] and conduct the additional experiments on SUN3D dataset [36] to further illustrate Key-Grid capability in coping with real and large-scale dataset.
@@ 4.3 Robustness Analysis
In this section, we focus on evaluating the robustness of Key-Grid on the noisy and down-sampled point clouds of folded shirts. To obtain the noisy point clouds, we add Gaussian noise with varying variances to the point clouds. And we utilize the Farthest Point Sampling method to downsample the original point clouds, which has also been used in previous works [17; 45].

### v2

@@ 1 Introduction
Representing objects through a set of 3D keypoints is one of the most popular and intuitive approaches to compress and comprehend 3D objects [31; 26]. Effectively exhibiting their utility, 3D keypoints have contributed to the success of a number of downstream tasks, including pose estimation, shape registration, object tracking in computer vision [15; 27; 4; 8; 44; 28; 34], as well as various kinds of robotic manipulation tasks [40; 1; 16].
@@
To make the detected keypoints as capable and accessible as possible, the research community is now concentrating on the unsupervised learning of semantically consistent keypoints on 3D point clouds. The implication of semantic consistency is typically twofold: the keypoints should be located at the semantically salient parts of the objects; they are also desired to be aligned within the same category even under large shape variations among diverse 3D object instances. To achieve these objectives, previous works [26; 9; 41; 3; 46] often adopt autoencoder frameworks to facilitate self-supervised training, where the encoders serving as the keypoint predictor are backbone networks [21; 22] generalizable to the shape variation, and the decoders reconstruct the input shape conditioned on the predicted keypoints. Since neural networks are in general better at compression than generation [20], the primary challenge of this pipeline lies in reconstructing the entire point cloud from a few estimated keypoints. Thus, the state-of-the-art (SOTA) detectors [26; 9] lay emphasis on leveraging different priors on 3D structures (e.g., “skeleton” in Skeleton Merger [26], and “cage” in KeypointDeformer [9]) so that the 3D object shapes can be more reasonably approximated by the information from the detected keypoints alone.
While maintaining semantic consistency is already demanding under the shape variations of rigid-body objects (e.g., those in the ShapeNetCoreV2 [2] dataset), it would be even more challenging if deformable objects are also taken into consideration. For instance, when detecting keypoints on the trousers, if one detected keypoint is located at one of the trouser legs, then the keypoint is desired to follow the motion of the trouser leg in the process of the trousers being folded (shown in Figure 1). Note that the shape variation caused by deformation is so dramatic that even the outline of the object shape has been completely altered, indicating a shift in the spatial and geometric structure of the keypoints. Despite the difficulty, with a growing interest in deformable object manipulation [1; 24; 25] in robotics as well as the emergence of large-scale datasets for deformable objects [49; 6] in computer vision, it is of increasing significance to develop a keypoint detector that is equally effectively when faced with deformable objects.
@@ 2 Related works
There have been various approaches proposed for supervised estimation of 3D keypoints using manually annotated keypoints [35; 5; 14; 52; 13; 42]. In contrast to supervised methods, our approach is unsupervised, meaning it does not rely on manually annotated keypoints. Thus, we review methods that adopt an unsupervised approach for estimating 3D keypoints. USIP [12] is the first detector that identifies 3D keypoints in an unsupervised manner by minimizing the chamfer distance between detected keypoints in randomly transformed object pairs. Canonical Capsules [29] is a similar approach that feeds pairs of a randomly translated copies of the same object into a network to detect keypoints. Following USIP, SC3K [51] also uses a random rotation to create two transformed visions of objects and generate the corresponding keypoints by mapping the keypoints of each version back to the original object. Another way to estimate 3D keypoints is proposed by Chen et al. [3]. They encode the point cloud as a set of local feature and input it into a novel structure model to generate the possibility of keypoints. Recently, Fernandez et al. [11] proposed a novel method that distinguishes between Node branches and Pose and coefficient branches to find the optimal keypoints of an object.
@@ 3.3 Decoder: Point Could Reconstruction
As an inverse process of the encoder, the decoder tries to reconstruct the entire point cloud by gradually augmenting finer geometric details in a hierarchical manner. Unlike previous methods [26; 9] where the keypoint-related 3D structures are used only once as the input of the encoder, we propose to integrate increasingly finer grid heatmap features with the layers of the decoder. In addition, the features in the encoder are directly copied and fused into the corresponding layer in the decoder in a U-Net [23] fashion to further assist the reconstruction process.
@@ 3.4 Training Objectives
The common way to evaluate the similarity between a reconstructed point cloud $\mathbf{X}_{r}$ and a target point cloud $\mathbf{X}$ is by employing the Chamfer distance metric [9; 48]. By minimizing the Chamfer distance between them, we can optimize the reconstruction process to closely match the target point cloud, thereby achieving a higher level of similarity between the two point clouds. Thus, we utilize the Chamfer distance between the reconstructed point cloud $\mathbf{X}_{r}$ and the target point cloud $\mathbf{X}$ to approximate this similarity loss $\mathcal{L}_{sim}$.
@@ 4.1 Setup
We use the ShapeNetCoreV2 and the ClothesNet datasets [2; 49] to evaluate the performance of Key-Grid. For the ShapeNetCoreV2 dataset [2], it contains 51,300 rigid-body objects of 55 different categories. In this paper, we only use categories that are manually annotated with semantic correspondence labels in the KeypointNet dataset [42]. For the ClothesNet dataset [49], we take three types of deformations on different type of clothing objects: dropping, dragging, and folding. For each deformation of different garment categories, there are 128 samples during the deformation process. We test the performance of Key-Grid on real-world 3D scans of clothing from the Deep Fashion3D V2 dataset [6] and conduct the additional experiments on SUN3D dataset [38] to further illustrate Key-Grid capability in coping with real and large-scale dataset.
@@ 4.3 Robustness Analysis
In this section, we focus on evaluating the robustness of Key-Grid on the noisy and down-sampled point clouds of folded shirts. To obtain the noisy point clouds, we add Gaussian noise with varying variances to the point clouds. And we utilize the Farthest Point Sampling method to downsample the original point clouds, which has also been used in previous works [19; 47].
@@ 5 Conclusion
Supplementary Materials
Appendix A Implementation Details
We use the PyTorch framework [7] to train our method. The entire network is optimized on the Adam optimizer [10] with a learning rate of 0.1. For all the experiments, the batch size is set to 8. Our method is trained for 100 epochs, and for the first 20 epochs of training, we do not use a similarity loss function which measures the similarity between the reconstructed point clouds and the original point clouds.
For all the experiments, the input point cloud contains $N=2048$ points. The number of encoder layers $L$ in the PointNet++ network [22] is 4. For the ShapeNetCoreV2 dataset [2], we set the number of keypoints $K$ to be 10. Thus, when calculating the farthest point keypoint loss function, we choose the number of the farthest points $J$ as 14. For the ClothesNet [49] and Deep Fash 3D dataset [6], we set $K$ and $J$ as 8 and 12, respectively. The hyperparameter $K$ and $J$ are 20 and 24 in the SUN3D dataset [38]. In the process of building the grid heatmap, the number of points $M$ uniformly selected on each edge of the cubic space is 16. For the decoder, the hyperparameter $N_{\mathrm{neig}}$ for the number of neighboring features applied in aligning the feature from the former layer to the current layer is set to 3. In the loss function, $\alpha_{far}$ and $\alpha_{sim}$ are both set to 1.
Appendix B Efficiency of Key-Grid
Table 6 presents the time and memory consumption of Key-Grid and SM when inferring a batch comprising 32 samples on a single 1080 Ti GPU. According to Table 6, Key-Grid achieves equal efficiency to the baseline method (SM) in terms of inference speed but consumes an acceptably larger memory size.
Method Memory (MiB) Time (s)
SM 4531 0.6
Ours 6081 0.6
Table 6: Analysis of Key-Grid efficiency.
Appendix C Visualization Results
In this section, we primarily present additional visualization results about the experimental part to illustrate the effectiveness of Key-Grid.
Number of keypoints.
We vary the number of unsupervised keypoints discovered by Key-Grid on the folded pants. Figure 7 shows that these different numbers of keypoints on the folded pants maintain good semantic consistency during the deformation process. And as the number of keypoints increases, we find that these keypoints do not overlap with each other, they still evenly distribute on the surface of the objects, and each keypoint represents the geometric information of the objects.
Figure 7: Different number of detected keypoints on the folded pants. Key-Grid identifies different numbers (6, 8, 10, 12) of keypoints on the folded pants, respectively.
Ablation study.
We show the visualization results of the ablation study in Figure 8. From the Figure 8, we find that in the decoder section, if we only use the encoder information or the grid heatmap information to reconstruct the original point cloud, the positions of the keypoints on the folded pants are not stable. When we do not use the similarity loss function in the decoder section, the keypoints will be placed in the wrong prior positions and the network predicts the redundant keypoints. Moreover, when we do not apply the farthest point keypoint loss function to train the network, the performance of keypoint recognition will be poor, as the keypoints will cluster around the center of the object.
Figure 8: Visualization results of the ablation study on the folded pants.
Gaussian noise to point cloud.
Figure 9 shows the visualization results of Key-Grid for different noisy point clouds, which add the Gaussian noise of different scales to the input point cloud. We can observe that even when Gaussian noise is added to the original point cloud with a magnitude of 0.08, the positions of the keypoints detected by Key-Grid in these noisy point clouds are still close to those found in the original point cloud. Thus, we can conclude that Key-Grid exhibits robustness when dealing with the noisy point clouds.
Figure 9: Performance of Key-Grid on the noisy point clouds. We indicate the level of Gaussian noise added underneath each noisy point cloud. “Noise 0.00” is the original point cloud.
Downsampling point cloud.
This paragraph presents the performance of Key-Grid on the downsampled point clouds shown in Figure 10. For decimating the original point cloud, we use the Farthest Point Sampling method for downsampling. We discover that Key-Grid successfully estimate the keypoints on the downsampled point clouds. When downsampling the original point cloud by a factor of 16, resulting in a point cloud with 128 sample points, except for the blue keypoints which will be shifted downwards, the positions of the other keypoints will remain close to their positions detected in the original point cloud.
Figure 10: Performance of Key-Grid on the downsampled point clouds. The input point clouds are downsampled for different scales, as mentioned at the bottom of each point cloud. “2048 sampled points” is the input point cloud.
Figure 11: Additional visualization results on the Deep Fash3D V2 dataset. Key-Grid recognizes the keypoints in these three garments with different deformation.
Deep Fash3D V2 dataset.
We show the additional visualization results of keypoints detected by Key-Grid on the Deep Fash3D V2 dataset [6]. Figure 13 illustrates that when faced with a wider variety of clothing cuff deformations, the keypoints detected by Key-Grid achieves great semantic consistency on the world scanned objects and represents the important information about these objects. Thus, we think that Key-Grid performs well in the world scanned objects.
Figure 12: Keypoint detection on the ShapeNetCoreV2 dataset. The keypoints are detected by SM and Key-Grid on the “Bed”, “Guitar”, “Car”, and “Motorcycle” category of objects in the ShapeNetCoreV2 dataset. Each category shows four samples.
Figure 13: Keypoint detection on the ClothesNet dataset. For the drop and drag deformation processes, we select the “Long Dress”, “Tie”, and “Vest” categories of objects to visualize the keypoints identified by SM and Key-Grid. We use lines to connect the keypoints at the same position during the deformation process.
ShapeNetCoreV2 and ClothesNet dataset.
For ShapeNetCoreV2 [2] and ClothesNet [49] dataset, we present the visualization results of keypoint recognition in a wider range of object categories using SM and Key-Grid. In Figure 12, we separately present the visualization results of keypoint recognition on the ShapeNetCoreV2 dataset using SM and Key-Grid. Compared to the redundancy phenomenon that many keypoints locate in the same region generated by SM [26], the keypoints detected by Key-Grid distribute separately on the surface of the object, and these keypoints also can represent the important geometric information of the object. For the ClothesNet dataset, Figure 13 demonstrates the keypoint detection of SM [26] and Key-Grid in the deformation process of a wider range of object categories, including drop and drag deformations. In the deformation process, compared to SM [26], the keypoints identified by our method keep good semantic consistency and distribute evenly on the surface of the object. Thus, for the rigid-body and deformable objects, Key-Grid exhibits great performance on keypoint detection.
Appendix D Video Results
In the supplementary materials, we also provide keypoint recognition results for the folding clothes in the ClothesNet dataset using different methods. The video we provide include the entire process of deformation for shirts and pants. From this video, we can conclude that Key-Grid identifies the keypoints that maintain good semantic consistency during the deformation process. These keypoints also carry the geometric feature information of the deformable objects.
Appendix E Limitation and Social Imapct
Following previous works, we need to manually set the total number of predicted keypoints beforehand. In future research, we will propose an adaptive keypoint detection method that can generate varying numbers of keypoints with accurate positions for different samples. Key-Grid predicts the keypoints on the publicly available datasets. Thus, we think Key-Grid has very limited potential negative societal impacts.